Grade 12th passAlgebra

Express each of the following as a product of prime factors only in exponential form:
108 × 192

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

The factors of 108 = 2 × 2 × 3 × 3 × 3
= 22 × 33
The factors of 192 = 2 × 2 × 2 × 2 × 2 × 2 × 3
= 26 × 3
Then,
=(22 ×33) × (26 x 3)
= 22 + 6× 33 +3 ... [∵am × an = am+n]
= 28 × 36

​Thanks